OpenStudy (anonymous):

find the range of the quadratic function. f(x) = (x+8)^2 - 7

OpenStudy (anonymous):

I think it is (-infinity, -7)

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

do i factor it out and find the maximum

OpenStudy (anonymous):

Put f(x) = y first: \[y = (x+8)^2 - 7\] \[(x+8)^2 = y + 7\] \[x + 8 = \sqrt{y + 7} \implies x = \sqrt{y + 7} - 8\]

OpenStudy (anonymous):

What values y cannot take??

OpenStudy (anonymous):

Number inside the square root must not be negative.. So: \[y + 7 \ge 0\] Can you solve now??

OpenStudy (anonymous):

Subtract both the sides by 7 and tell me what do you get??

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

so is it (-7, infinity)

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

thank you again